What is the length of FH If G is the midpoint, FG =12x-5,and GH = 7x?
Answer by jim_thompson5910(35256)

G is the midpoint, so the two smaller pieces are equal which means FG = GH

FG = GH

12x - 5 = 7x

- 5 = 7x-12x


- 5 = -5x


-5x = -5

x = -5/(-5)

x = 1

-------------------------------------------------------

FH = FG + GH

FH = (12x - 5) + (7x)

FH = (12*1 - 5) + (7*1)

FH = (12 - 5) + (7)

FH = 12 - 5 + 7

FH = 7 + 7

FH = 14